OpenStudy (anonymous):

Which is equivalent to 3-2 ( x + 7 ) = 12? A. 3-2x - 14 = 12 b. 3-2x + 7 = 12 C. -2x - 14 = 15 D.x - 14 = 12

OpenStudy (anonymous):

@Hanishala

OpenStudy (anonymous):

option A...simply multiply the terms inside racket.. @keepsecure93

OpenStudy (anonymous):

it will be 3 -2x -14 =12 yes it will be A

OpenStudy (anonymous):

which equation is equivalent to 2/4 ( 3x + 9) = 21 A. 2x + 9 = 21 B. 6x + 18 = 21 C. 3x + 6 = 21 D. 2x + 6 = 21

OpenStudy (anonymous):

@Hanishala @zaibali.qasmi

OpenStudy (anonymous):

it should be 3x+6=42 or if it is 4/2 then 6x+18=21 see it again carefully..

OpenStudy (anonymous):

The number line below shows the solution set for which of the following equations?